a. Analyse the cross price elasticity of demand for wheat and rice when a change in the price of wheat from Rs.70 to Rs. 90 results in a change in the quantity demand for wheat from kg 3000 to kg 5000 in the market. Interpret the value of the coefficient.


1
Expert's answer
2021-04-20T12:47:10-0400

Formula for the cross price elasticity of demand is : 

Ed=2(Q2Q1)(Q2+Q1)×(P2+P1)2(P2P1)Ed = \frac{2(Q2-Q1)}{(Q2+Q1)} × \frac{(P2+P1)}{2(P2-P1)}

Ed=2(50003000)(5000+3000)×(90+70)2(9070)Ed =\frac {2 (5000-3000)}{(5000+3000)} ×\frac{ (90+70)}{2(90-70)}  


Ed=40008000×16040Ed = \frac{4000}{8000} × \frac{160}{40}


Ed=12×4Ed = \frac{1}{2} × 4


Ed=2.Ed = 2.


The cross price elasticity of demand is equal to 2 and it is greater than 1 and therefore the demand is elastic and it is positive because of the wheat and rice are complementary goods and with the increase in the price of the wheat the demand for the rice increases.